OpenStudy (anonymous):

Does anyone know how to factor trinomials completely? I need to factor: 20a - 2a^ - 18 completely. Here is what I have so far: 2(10A - a^ -9). What do I do next? The solution is: -2(a - 9)(a - 1) but I cannot figure out how to get there.

OpenStudy (anonymous):

you need to put the a^2 term first

OpenStudy (anonymous):

Use reverse FOIL?

OpenStudy (anonymous):

I understand I need to put the a ^ 2 first, but I don't understand WHY I need too?? How do you know to put the a ^2 first?

OpenStudy (anonymous):

-2a^2 + 20a - 18 factor out a -2 -2(a^2 - 10a + 9) you need factors of 9 that add to get 10 ---- 9 and 1 -2(a - 9)(a - 1)

OpenStudy (anonymous):

-2(a^2-10a+9)=0 (a-9)(a-1)=0 a=1,9

OpenStudy (anonymous):

the form of a quadratic is ax2+bx+c

OpenStudy (anonymous):

great ankur!

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

Think about... what 2 numbers times each other equals -9 and added together equals -10.. Also always make sure the a^2 term is positive so it will make the factoring very easy